### v1

@@ Abstract
Context. The Milky Way’s Central Molecular Zone (CMZ) is measured to form stars 10 times less efficiently than in the Galactic disk, based on emission from high-mass stars. However, the CMZ’s low-mass (2 ) protostellar population, which accounts for most of the initial stellar mass budget and star formation rate (SFR), is poorly constrained observationally due to limited sensitivity and resolution.
@@
Methods. We present the Dual-band Unified Exploration of Three CMZ Clouds (DUET) survey, targeting the the 20 km s-1 cloud, Sgr C, and Dust Ridge cloud e using the the Atacama Large Millimeter/submillimeter Array (ALMA) at 1.3 and 3 mm. The mosaicked observations achieve a comparable resolution of 0.2–0.3″ (1600–2500 au) and a sky coverage of 8.3–10.4 arcmin2, respectively.
Results. We report 563 continuum sources at 1.3 mm and 330 at 3 mm, respectively, and a dual-band catalog with 450 continuum sources. These sources are marginally resolved at the 2,000 au resolution. We find a universal deviation (across 70% of the source sample) from commonly-used dust modified blackbody (MBB) models, characterized by either low spectral indices or low brightness temperatures.
Conclusions. Three possible explanations are discussed for the deviation. (1) Optically thick Class 0/I Young stellar objects (YSOs) with very small beam filling factor can lead to lower brightness temperatures than what MBB models predict. (2) Large (mm/cm-sized) dust grains have more significant self-scattering, and therefore frequency-dependent albedo could cause lower spectral indices. (3) Free-free emission over 30 Jy can severely contaminate dust emission and cause low spectral indices for mJy sources in our sample, although the needed number of massive protostars (embedded UCHii regions) are infeasibly high for the normal stellar initial mass function. A reliable measurement of the SFR at low protostellar masses will require future work to distinguish between these possible explanations.
@@
The Central Molecular Zone (CMZ), the inner 500 pc of the Milky Way, contains more than of dense molecular gas (Morris & Serabyn 1996; Ferrière et al. 2007). It is a unique star-forming environment, characterized by extreme conditions similar to those under which most of the cosmic star formation may have occurred (Mezger et al. 1996; Kruijssen & Longmore 2013). These conditions include strong magnetic fields, intense turbulence, elevated cosmic ray flux, and gravitational shear, along with feedback from the supermassive black hole Sgr A⋆ (Genzel & Townes 1987; Kruijssen et al. 2014; Mills 2017; Bryant & Krabbe 2021; Henshaw et al. 2023). Several molecular clouds with masses over and densities of 104 cm-3 are found within the CMZ (Bally et al. 1987, 1988; Lis & Carlstrom 1994; Tsuboi et al. 1999; Kendrew et al. 2013; Walker et al. 2015; Kauffmann et al. 2017a, b; Lu et al. 2019b, 2020; Battersby et al. 2020). Despite the presence of a large reservoir of molecular gas at volume densities high enough to generate widespread star formation in the solar neighborhood (e.g. Lada et al. 2010), the CMZ is notoriously inefficient at forming stars, with a star formation rate (SFR) of yr-1 inferred from massive ( ) stellar emission (see review Table. 1 in Henshaw et al. 2023) that is about 10 times lower than expected based on dense gas star formation relations (e.g. Longmore et al. 2013; Barnes et al. 2017).
@@
The measurements of the SFR in the CMZ initially relied on free-free emission from massive stars at extremely low angular resolution (Murray & Rahman 2010). Follow-up interferometric observations have provided direct evidence of high-mass ( ) star formation. For instance, observations with the Submillimeter Array (SMA) have resolved several hundreds of dense cores at 0.2 pc resolution (Liu et al. 2013; Kauffmann et al. 2017a; Lu et al. 2015, 2017, 2019b; Hatchfield et al. 2020). The Very Large Array (VLA) surveys found Class ii CH3OH masers, H2O masers, and ultra-compact Hii (UCHii) regions associated with these dense cores of – (Lu et al. 2019a, b).
@@
However, the low-mass ( ) protostellar population in the CMZ remains unclear limited by resolution and sensitivity. Theories suggest that both high-mass and low-mass protostars form in a common accretion reservoir; therefore, low-mass protostars should be found around high-mass ones (e.g., Bonnell et al. 2001; Smith et al. 2009). Recent high-resolution and high-sensitivity Atacama Large Millimeter/submillimeter Array (ALMA) observations toward star-forming regions in the Galactic disk detect faint low-mass cores surrounding massive ones, revealing crowded mini-clusters that were previously identified as massive cores (e.g. Wang et al. 2014; Zhang et al. 2015; Pillai et al. 2019; Xu et al. 2023, 2024b). These low-mass members are crucial for measuring the SFR. For a normal stellar initial mass function (IMF, e.g. Bastian et al. 2010), they represent the dominant source of both the stellar mass and the star formation rate. In particular, a top-heavy IMF has been suggested for young massive clusters in the CMZ (Figer et al. 1999; Wang et al. 2006; Hosek et al. 2019), and studying low-mass star formation in CMZ protoclusters will be crucial to place these claims in the context of the incipient stellar population.
2 DUET: ALMA dual-band mosaic of CMZ clouds
@@
To investigate low-mass star formation in the CMZ, we have conducted the Dual-band Unified Exploration of Three CMZ Clouds (DUET) observing project at both 1.3 mm and 3 mm by ALMA. Another feature of DUET is that the fields of view of are as large as 12–37 at 1.3 mm and 18–64 at 3 mm by mosaicking ALMA pointings. DUET serves as part of the Coordinated Observations of Nebulae in the CMZ Exploring Gas Recycling and Transformation (CONCERT) campaign. To achieve sufficient detections of low-mass fragments, we select three massive molecular clouds namely the 20 km s-1 cloud, Sgr C, and cloud e, which present abundant fragmentation in the 1.3 mm continuum emission (Lu et al. 2020). The three clouds, together with a few other clouds in the CMZ, including G0.253+0.016 and the 50 km s-1 cloud, were originally in a mini-survey with the SMA and VLA to study their star formation and dense core population (Kauffmann et al. 2017a, b). The sample covers all major molecular clouds in the CMZ outsider of Sgr B2. Compared to G0.253+0.016 and the 50 km s-1 cloud, the three DUET clouds have been found to contain a larger number of dense cores (Rathborne et al. 2015; Lu et al. 2019b; Walker et al. 2021; Lu et al. 2020), and are therefore selected for DUET sample.
In the Galactic context, the 20 km s-1 cloud (M0.130.08) is projected to within 10 pc of Sgr A⋆. It is suggested to feed gas and dust to the 2-pc circum-nuclear disk (Ho et al. 1991), and shows a moderate level of star formation itself (Kauffmann et al. 2017a; Lu et al. 2019b, 2021). Sgr C is the most massive and luminous star-forming region on the western side of the CMZ, and it has been suggested to be a connection point to a stream of gas and dust linking the CMZ and the Nuclear Stellar Disk to the Galactic bar (Henshaw et al. 2023). Star formation in Sgr C has been found to be active, with an overall efficiency comparable to those in Galactic disk clouds (Kauffmann et al. 2017a; Lu et al. 2019b). Dust Ridge cloud e is one of the massive clouds ( ) along the Dust Ridge extending to the eastern side of Sgr A⋆ (Immer et al. 2012). It is largely devoid of signatures of ongoing star formation, except for two dense cores associated with masers and a UCHii region candidate (Lu et al. 2019b, a). Therefore, the three clouds constitute a small yet representative sample in terms of their locations in the CMZ and their levels of star formation. Fig. 1 shows an overview of the three clouds, with their properties including the cloud radii, masses, and H2 volume densities as listed in Table 1.
@@
The ambiguous nature of millimeter emission in these clouds is our primary motivation. Previous ALMA single-band 1.3 mm continuum observations have revealed hundreds of compact sources on 2000-au scales, with flux densities as low as 1 mJy (Lu et al. 2020). Assuming the 1.3 mm emission arises purely from optically thin dust at a uniform temperature of 25 K, these compact continuum sources correspond to a population of low-mass dense cores with masses between 0.6 and 2 . However, there are several caveats in using a single-band millimeter continuum observation to characterize the dense cores. First, dust emission is not always optically thin, particularly at shorter wavelengths, potentially leading to an underestimation of the derived masses. Second, these sources may consist of a mix of thermal and non-thermal emissions, particularly considering the widespread non-thermal radio continuum emission associated with strong magnetic fields or cosmic rays in the CMZ (Yusef-Zadeh et al. 2013; Heywood et al. 2022; Bally et al. 2024). Third, thermal free-free emission from ionized gas can contaminate the millimeter continuum (e.g., Ginsburg et al. 2018; Lu et al. 2019b). The dual-band feature of DUET allows us to disentangle these complications, providing crucial insights into the nature of these compact sources.
In the following, we introduce the ALMA observations and data reduction in Sect. 3. We present the source extraction processes, the derived source catalogs, and the quality assessment of the catalogs in Sect. 4. We then discuss the intrinsic source size in Sect. 5. Leveraging the dual-band catalog, we find a universal ( 70% in the source sample) deviation from commonly-used dust modified blackbody emission models, and discuss three possible interpretations in Sect. 6: Class 0/I young stellar objects with very small source size, existence of large dust grains, and free-free emission contamination. Throughout this paper, we adopt a uniform heliocentric distance of 8.277 kpc to these three clouds (GRAVITY Collaboration et al. 2022).
@@ 3.1 ALMA 3 mm
The correlators were configured with four spectral windows to cover several important spectral lines. The frequency coverage and an example of observed spectra are presented in Fig. 11. Two 0.9375-GHz spectral windows with a spectral resolution of 0.488 MHz (1.67 km s-1) were put at central rest frequencies of 86.71 GHz and 88.29 GHz, respectively. Two 1.875-GHz spectral windows with a spectral resolution of 0.976 MHz (2.95 km s-1) were centered at the rest frequencies of 98.10 GHz and 100.00 GHz, respectively. They were set to cover dense gas tracers including H13CN (1–0), H13CO+ (1–0), HN13C (1–0), HCN (1–0), OCS (8–7), and CS (2–1), shock tracers including SiO (2–1), SO v=0 3(2)–2(1), CH3OH 2(1,1)–1(1,0)A, and HNCO 4(0,4)–3(0,3), and an ionized gas tracer H40. We mark several strong lines in Fig. 11, but note that there are plenty of complex organic molecule (COM) lines unlabeled. Analyses of COM lines will be presented in forthcoming papers of the series. A total bandwidth of 5.6 GHz for the continuum was achieved, although when we imaged the continuum, a portion of the bandwidth was flagged to remove spectral lines.
@@
All the data were calibrated using the standard ALMA pipeline (Hunter et al. 2023) implemented in the Common Astronomy Software Applications package (CASA; CASA Team et al. 2022). Due to the three different epochs of observations, three different CASA versions were used, including 5.4.0-68, 6.1.1-15, and 6.2.1-7. Channels with spectral line emission in the visibility data were visually identified with the plotms task in CASA, and were flagged to keep only line-free channels. For 20 km/s cloud, there is 6.3% of the bandwidth flagged. This is 4.1% for Sgr C and 9.2% for cloud e. The calibrated and flagged visibility data of the C43-4 and C43-7 configurations were then merged using the concat task.
@@
Continuum images were reconstructed from the line-free channels, using the multi-term, multi-frequency synthesis method (mtmfs) with nterm=2 in the tclean task in CASA 6.5.2-26. Briggs weighting with a robust parameter of 0.5 was used. The multi-scale deconvolution algorithm was adopted with multi-scale parameters of [0, 5, 15, 50, 150] for a pixel size of 004. The reference frequency of the continuum map produced by mtmfs is 93.6 GHz. For the 20 km s-1 cloud and Sgr C where multiple pointings were observed, we produced the mosaicked images. The resulting synthesized beam sizes are on average 032022 (equivalent to 26001800 au). Due to the lack of short baselines, the data are not sensitive to structures larger than 14″ (0.56 pc). This value is larger than that of 1.3 mm by a factor of , which potentially bias the spectral index analyses. So, when discussing spectral index in Sect. 6.1, we will re-clean the continuum image by constraining the same -range at two bands. The 3 mm continuum image is published in 10.5281/zenodo.15068333, and its emission is shown in red color in Figs. 2, as compared to the cyan color showing the 1.3 mm emission. The continuum rms noise in Col. 4 of Table 2 was measured on the image without primary beam corrections with uniform thermal noise by iteratively clipping until convergence at around its median (Stetson 1987; Bertin & Arnouts 1996).
@@
The ALMA 1.3 mm continuum and spectral line data have been reported in Lu et al. (2020, 2021). Details of the observation setups and data calibration can be found therein. The ALMA project code is 2016.1.00243.S. We highlight relevant information as follows. We have used the standard CASA pipeline to calibrate the data, and used the tclean task in CASA to image the visibility data after combining data from two array configurations C40-5 and C40-3. The continuum reference frequency is 226.0 GHz. The properties of the Band-6 continuum images are summarized in Sect. 2. The pixel size is 004, the same as that of the Band 3 images. The resulting synthesized beam is 025017 (equivalent to 2000 au 1400 au). The maximum recoverable scale is 64 (0.26 pc). The measured continuum noise rms is 43–45 Jy beam-1. The 1.3 mm continuum image is published in 10.5281/zenodo.15068333, and its emission is shown in cyan color in Figs. 2, as compared to the red color showing the 3 mm emission.

### v2

@@ Abstract
Context. The Milky Way’s central molecular zone (CMZ) has been measured to form stars ten times less efficiently than in the Galactic disk, based on emission from high-mass stars. However, the CMZ’s low-mass (2 ) protostellar population, which accounts for most of the initial stellar mass budget and star formation rate (SFR), is poorly constrained observationally due to limited sensitivity and resolution.
@@
Methods. We present the Dual-band Unified Exploration of three CMZ Clouds (DUET) survey, targeting the the 20 km s-1 cloud, Sgr C, and the dust ridge cloud “e” using the Atacama Large Millimeter/submillimeter Array (ALMA) at 1.3 and 3 mm. The mosaicked observations achieve a comparable resolution of 02–03 (2000 au) and a sky coverage of 8.3–10.4 arcmin2, respectively.
Results. We report 563 continuum sources at 1.3 mm and 330 at 3 mm, respectively, and a dual-band catalog with 450 continuum sources. These sources are marginally resolved at a resolution of 2000 au. We find a universal deviation (% of the source sample) from commonly used dust modified blackbody (MBB) models, characterized by either low spectral indices or low brightness temperatures.
Conclusions. Three possible explanations are discussed for the deviation. (1) Optically thick class 0/I young stellar objects (YSOs) with a very small beam filling factor can lead to lower brightness temperatures than what MBB models predict. (2) Large dust grains with millimeter or centimeter in size have more significant self-scattering, and frequency-dependent albedo could therefore cause lower spectral indices. (3) Free-free emission over 30 Jy can severely contaminate dust emission and cause low spectral indices for milli-Jansky sources, although the number of massive protostars (embedded UCHii regions) needed is infeasibly high for the normal stellar initial mass function. A reliable measurement of the SFR at low protostellar masses will require future work to distinguish between these possible explanations.
@@
The central molecular zone (CMZ), the inner pc of the Milky Way, contains more than of dense molecular gas (Morris & Serabyn 1996; Ferrière et al. 2007). It is a unique star-forming environment, characterized by extreme conditions similar to the ones under which most of the cosmic star formation may have occurred (Mezger et al. 1996; Kruijssen & Longmore 2013). These conditions include strong magnetic fields, intense turbulence, elevated cosmic ray flux, and gravitational shear, along with feedback from the supermassive black hole Sgr A⋆ (Genzel & Townes 1987; Kruijssen et al. 2014; Mills 2017; Bryant & Krabbe 2021; Henshaw et al. 2023). Several molecular clouds with masses of over and densities of 104 cm-3 are found within the CMZ (Bally et al. 1987, 1988; Lis & Carlstrom 1994; Tsuboi et al. 1999; Kendrew et al. 2013; Walker et al. 2015; Kauffmann et al. 2017a, b; Lu et al. 2019b, 2020; Battersby et al. 2020). Despite the presence of a large reservoir of molecular gas at volume densities high enough to generate widespread star formation in the solar neighborhood (e.g., Lada et al. 2010), the CMZ is notoriously inefficient at forming stars, with a star formation rate (SFR) of yr-1 inferred from massive ( ) stellar emission (see review Table. 1 in Henshaw et al. 2023) which is about ten times lower than expected based on dense gas star formation relations (e.g., Longmore et al. 2013; Barnes et al. 2017).
@@
The measurements of the SFR in the CMZ initially relied on free-free emission from massive stars at an extremely low angular resolution (Murray & Rahman 2010). Follow-up interferometric observations have provided direct evidence of high-mass ( ) star formation. For instance, observations with the Submillimeter Array (SMA) have resolved several hundred dense cores at 0.2 pc resolution (Liu et al. 2013; Kauffmann et al. 2017a; Lu et al. 2015, 2017, 2019b; Hatchfield et al. 2020). The Very Large Array (VLA) surveys found class ii CH3OH masers, H2O masers, and ultracompact Hii (UCHii) regions associated with these dense cores of – (Lu et al. 2019a, b).
@@
However, the low-mass ( ) protostellar population in the CMZ remains unclear, limited by resolution and sensitivity. Theories suggest that both high-mass and low-mass protostars form in a common accretion reservoir. Therefore, low-mass protostars should be found around high-mass ones (e.g., Bonnell et al. 2001; Smith et al. 2009). Recent high-resolution and high-sensitivity Atacama Large Millimeter/submillimeter Array (ALMA) observations toward star-forming regions in the Galactic disk have detected faint low-mass cores surrounding massive ones, revealing crowded mini-clusters that were previously identified as massive cores (e.g., Wang et al. 2014; Zhang et al. 2015; Pillai et al. 2019; Xu et al. 2023, 2024b). These low-mass members are crucial to measuring the SFR. For a normal stellar initial mass function (IMF, e.g., Bastian et al. 2010), they represent the dominant source of both the stellar mass and the SFR. In particular, a top-heavy IMF has been suggested for young massive clusters in the CMZ (Figer et al. 1999; Wang et al. 2006; Hosek et al. 2019), and studying low-mass star formation in CMZ protoclusters will be crucial to place these claims in the context of the incipient stellar population.
2 DUET: dual-band mosaic of three CMZ clouds
@@
To shed light on the low-mass population in the Milky Way CMZ, we have conducted the Dual-band Unified Exploration of three CMZ Clouds (DUET) survey at both 1.3 mm and 3 mm by ALMA. DUET serves as part of the Coordinated Observations of Nebulae in the CMZ Exploring Gas Recycling and Transformation (CONCERT) campaign.
Our sample selection took advantage of the previous SMA survey of all major molecular clouds in the CMZ outsider of Sgr B2 (Kauffmann et al. 2017a, b). We selected three massive molecular clouds namely the 20 km s-1 cloud, Sgr C, and cloud e, which present abundant fragmentation in the 1.3 mm continuum emission (Lu et al. 2020). Compared to G0.253+0.016 and the 50 km s-1 cloud (Rathborne et al. 2015; Walker et al. 2021), the three DUET clouds have been found to contain a larger number of dense cores.
In the Galactic context, the 20 km s-1 cloud (M0.130.08) is projected to within 10 pc of Sgr A⋆. It has been suggested that it feeds gas and dust to the 2-pc circum-nuclear disk (Ho et al. 1991), and it shows a moderate level of star formation itself (Kauffmann et al. 2017a; Lu et al. 2019b, 2021). Sgr C is the most massive and luminous star-forming region on the western side of the CMZ, and it has been suggested that it is a connection point to a stream of gas and dust linking the CMZ and the nuclear stellar disk to the Galactic bar (Henshaw et al. 2023). Star formation in Sgr C has been found to be active, with an overall efficiency comparable to ones in Galactic disk clouds (Kauffmann et al. 2017a; Lu et al. 2019b). The dust ridge cloud “e” is one of the massive clouds ( ) along the dust ridge extending to the eastern side of Sgr A⋆ (Immer et al. 2012). It is largely devoid of signatures of ongoing star formation, except for two dense cores associated with masers and a UCHii region candidate (Lu et al. 2019b, a). Therefore, the three clouds constitute a small yet representative sample in terms of their locations in the CMZ and their levels of star formation. Fig. 1 shows an overview of the three clouds, with their properties including the cloud radii, masses, and H2 volume densities listed in Table 1.
@@
DUET primary motivation lies in penetrating the ambiguous nature of millimeter emission in these clouds. Previous ALMA single-band 1.3 mm continuum observations have revealed hundreds of compact sources on 2000-au scales, with flux densities as low as 1 mJy (Lu et al. 2020). Assuming that the 1.3 mm emission arises purely from optically thin dust at a uniform temperature of 25 K, these compact continuum sources correspond to a population of low-mass dense cores with masses between 0.6 and 2 . However, there are several caveats in using a single-band millimeter continuum observation to characterize the dense cores. First, dust emission is not always optically thin, particularly at shorter wavelengths, potentially leading to an underestimation of the derived masses. Second, these sources may consist of a mix of thermal and non-thermal emissions, particularly considering the widespread non-thermal radio continuum emission associated with strong magnetic fields or cosmic rays in the CMZ (Yusef-Zadeh et al. 2013; Heywood et al. 2022; Bally et al. 2024). Third, thermal free-free emission from ionized gas can contaminate the millimeter continuum (e.g., Ginsburg et al. 2018; Lu et al. 2019b). Last but not least, previous observations are mostly single pointings, which could miss the population outside fields. The dual-band mosaics by DUET allow us to address these problems, providing an unprecedented insights into the nature of these cloud-wide compact sources.
In the following, we introduce the ALMA observations and data reduction in Sect. 3. We present the source extraction processes, the derived source catalogs, and the quality assessment of the catalogs in Sect. 4. We then discuss the intrinsic source size in Sect. 5. Leveraging the dual-band catalog, we find a universal ( 70% in the source sample) deviation from commonly used dust modified blackbody (MBB) emission models, and discuss three possible interpretations – class 0/I young stellar objects (YSOs) with a very small source size, the existence of large dust grains, and free-free emission contamination – in Sect. 6. Throughout this paper, we adopt a uniform heliocentric distance to these three clouds of 8.277 kpc (GRAVITY Collaboration et al. 2022).
@@ 3.1 ALMA 3 mm
The correlators were configured with four spectral windows to cover several important spectral lines. The frequency coverage and an example of observed spectra are presented in Fig. 11. Two 0.9375-GHz spectral windows with a spectral resolution of 0.488 MHz (1.67 km s-1) were put at central rest frequencies of 86.71 GHz and 88.29 GHz, respectively. Two 1.875-GHz spectral windows with a spectral resolution of 0.976 MHz (2.95 km s-1) were centered at the rest frequencies of 98.10 GHz and 100.00 GHz, respectively. They were set to cover dense gas tracers including H13CN (1–0), H13CO+ (1–0), HN13C (1–0), HCN (1–0), OCS (8–7), and CS (2–1), shock tracers including SiO (2–1), SO v=0 3(2)–2(1), CH3OH 2(1,1)–1(1,0)A, and HNCO 4(0,4)–3(0,3), and an ionized gas tracer, H40. We mark several strong lines in Fig. 11, but note that there are plenty of unlabeled complex organic molecule (COM) lines. Analyses of COM lines will be presented in forthcoming papers of the series. A total bandwidth of 5.6 GHz for the continuum was achieved, although when we imaged the continuum a portion of the bandwidth was flagged to remove spectral lines.
@@
All the data were calibrated using the standard ALMA pipeline (Hunter et al. 2023) implemented in the Common Astronomy Software Applications package (CASA; CASA Team et al. 2022). Due to the three different epochs of observations, three different CASA versions were used, including 5.4.0-68, 6.1.1-15, and 6.2.1-7. Channels with spectral line emission in the visibility data were visually identified with the plotms task in CASA, and were flagged to keep only line-free channels. For the 20 km/s cloud, 6.3% of the bandwidth is flagged. This is 4.1% for Sgr C and 9.2% for cloud e. The calibrated and flagged visibility data of the C43-4 and C43-7 configurations were then merged using the concat task.
@@
Continuum images were reconstructed from the line-free channels, using the multi-term, multifrequency synthesis method (mtmfs) with nterm=2 in the tclean task in CASA 6.5.2-26. Briggs weighting with a robust parameter of 0.5 was used. The multi-scale deconvolution algorithm was adopted with multi-scale parameters of [0, 5, 15, 50, 150] for a pixel size of 004. The reference frequency of the continuum map produced by mtmfs is 93.6 GHz. For the 20 km s-1 cloud and Sgr C, where multiple pointings were observed, we produced the mosaicked images. The resulting synthesized beam sizes are on average 032022 (equivalent to 26001800 au). Due to the lack of short baselines, the data are not sensitive to structures larger than 14″ (0.56 pc). This value is larger than that of 1.3 mm by a factor of , which potentially bias the spectral index analyses. So, when discussing spectral index in Sect. 6.1, we re-cleaned the continuum image by constraining the same -range at two bands. The 3 mm continuum image is published in 10.5281/zenodo.15068333, and its emission is shown in a red color in Fig. 2, compared to the cyan color showing the 1.3 mm emission. The continuum noise root mean square (rms) in Col. 4 of Table 2 was measured on the image without primary beam corrections with uniform thermal noise by iteratively clipping until convergence at around its median (Stetson 1987; Bertin & Arnouts 1996).
@@
The ALMA 1.3 mm continuum and spectral line data have been reported in Lu et al. (2020, 2021). Details of the observation setups and data calibration can be found therein. The ALMA project code is 2016.1.00243.S. We highlight relevant information as follows. We have used the standard CASA pipeline to calibrate the data, and used the tclean task in CASA to image the visibility data after combining data from two array configurations C40-5 and C40-3. The continuum reference frequency is 226.0 GHz. The properties of the band 6 continuum images are summarized in Sect. 2. The pixel size is 004, the same as that of the band 3 images. The resulting synthesized beam is 025017 (equivalent to 2000 au 1400 au). The maximum recoverable scale is 64 (0.26 pc). The measured continuum noise rms is 43–45 Jy beam-1. The 1.3 mm continuum image is published in 10.5281/zenodo.15068333, and its emission is shown in a cyan color in Fig. 2, compared to the red color showing the 3 mm emission.